Sol Summary – May 8th

Sol Summary – May 8th

Crew 212 Sol 3 Summary Report 8-MAY-2019

Sol: 3

Summary Title: Bad weather is hating our experiments part 1

Author’s name: Marlen Castillo Vilcahuaman

Mission Status: Things are getting heated up! But ok.

Sol Activity Summary: We had a morning EVA. Lots of lessons to learn after that EVA. In the morning, Commander fell under the spell of a cold sore. Lip creams are adviced. We had a soup thanks to our creative crew members. Right now, waiting for the bread!

Look Ahead Plan: 2 EVAs for tomorrow, depending on the weather and the information we can achieve tonight.

Anomalies in work: Commander and her cold sore.

Weather: Strong winds with rain in the afternoon.

Crew Physical Status: … Commander and her cold sore is all that needs to be reported.

EVA: 1

Reports to be file: Operations Report, GreenHab Report, Journalist Report, Daily Photo Report (not zipped), EVA Request, Space Suit Report, Astronomy report, Space suit report, HSO report

Support Requested: None.

Operations Report – May 8th

Crew 212 Operations Report 08-MAY-2019

SOL: 3

Name of person filing report: Paolo Guardabasso

Non-nominal systems: Oven in the kitchen

Notes on non-nominal systems: The kitchen oven door doesn’t fully close and it was necessary to block it with some tape to use it.

Generator (hours run last night): 12:30

On (last night): 6:30 PM

Off (this morning): 7:00 AM

Solar— SOC 86% at 7:40PM

Diesel Reading – 97%

Propane Reading – 46%

Ethanol Free Gasoline – 4 gallons

Water (Axillary tank) – 0 gallons

Water (static tank) – 270 gallons

Axillary to Static tank transfer – no

Gallons transferred: No

Water in GreenHab – 170 gallons

Water (loft) – Static to Loft Pump used – Yes

Water Meter: 1433682

Toilet tank emptied: Yes

Deimos rover used: no (out of order)


Beginning charge:

Ending charge:

Currently charging:

Sojourner rover used: ASSIGNED TO DIRECTOR


Beginning charge:

Ending charge:

Currently charging:

Spirit rover used: Yes

Hours: 92.1h (the previous value of 92.6h was wrong, the correct one was 91.6h)

Beginning charge: 100%

Ending charge: 80%

Currently charging: Yes

Opportunity rover used: No

Hours: 58.7h

Beginning charge:

Ending charge:

Currently charging: No

Curiosity rover used: Yes

Hours: 94.3h

Beginning charge: 100%

Ending charge: 95%

Currently charging: Yes

Notes on rovers: An Engineering EVA was done to recover a payload in Spirit’s trunk.

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3)

Reason for use:

Oil Added?

ATV Fuel Used: Gals

# Hours the ATVs were Used today:

Notes on ATVs:

HabCar used and why, where? Nothing to report

CrewCar used and why, where? Off Campus

General notes and comments: Today’s EVA was shortened because of the weather.

Summary of internet: Nothing to report

Summary of suits and radios: A major check was run, and each crew member was assigned a functioning radio.

Summary of Hab operations: The oven door doesn’t fully close, so we closed it with tape to cook biscuits. A leak on the lower floor was found after the rain at noon, close to the Engineering Airlock door (right above the propane line).

Summary of GreenHab operations: Nothing to report

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: It was depressurized to allow the

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing to report

Greenhab Report – May 8th

Crew 212 GreenHab Report – 8-May-2019: SOL 3

GreenHab Officer: Hector Palomeque

Environmental control: Ambient with door open from 07:00 to 11:00

Shade cloth 80%: on

Average temperatures:
Low: 15.9° C
High: 35° C

Hours of supplemental light: 0

Daily water usage for crops: 4 gallons

Water in Blue Tank: 169 gallons

Time(s) of watering for crops: 07:00

Changes to crops: All plants look healthy. Chard is ready to harvest. Basil is still blooming. Tomato in the blue tray is getting yellow in some leaves, might be hydric stress.

Narrative: Plants were watered just one time due to the weather that is cold and the pots/trays still with enough moisture.

Harvested: 353g Carrots (wet weight)

Support/supplies needed: None.

Astronomy Report – May 8th

Astronomy Report
Name: Mariona Badenas
Crew: 212
Date: 8 May 2019

Robotic Telescope Requested (choose one): MDRS-WF

Object Viewed #1: NGC4236
Problems Encountered: None (see attached picture)

Object Viewed #2: IC5067 (very excited about this one!)
Problems Encountered: The telescope was open last night, but Skynet reported a mount error. The observations are still not complete.

(More objects to be Viewed): NGC6992 and the Veil Nebula
Problems Encountered: N/A (Observations of these objects were requested this afternoon).

Solar Features Observed: None
Problems Encountered: Bad weather.

EVA Report – May 8th

Crew 212 EVA Report 08-MAY-2019

EVA #03

Author: Zoe Townsend

Purpose of EVA: First testing of drones, soils collection and thermal rover tests.

Start time: 09:00 AM

End time: 09:55 AM

Narrative: This was an EVA aimed at simultaneously running three experiments at once. However, at 07:00H we received a message from Shannon warning us about weather in the area and the potential of cancelling the EVA. It was decided at 08:30 that the weather was calm enough to attempt an EVA, however miscommunication between the crew meant that some believed the EVA was still 2hours whilst other believed it had been shortened to 1 hour from weather constraints. Due to the late start and the equipment required, the crew officially left the Hab at 09:30.

The first of the experiments was to collected soil, rock and sand samples. Whilst soil was collected, due to the location chosen and limited times the rock and sand were not collected in time in sufficient qualities and therefore the collection will be to be completed on another EVA.

The second experiment was to test the effectiveness of the thermal shielding that will be used as part of the rover. However, overconfidence with the flexibility of the suit meant that the experiment was planned for one person. Unfortunately, because of the wind and the fact the experiment takes place on the floor, two people are required next time to prevent damaging the blanket.

The final experiment was the first testing of the drones. The limited time frame prior to the mission meant that the RAM was not organised for pressurization. The contributed to the slow leaving of the HAB. When finally at the site the experiment was unable to be assembled in the timeframe.

Whilst none of the experiments were completed to satiation a number of lessons have me learnt and are summarised below from the after EVA debrief.

· Do a pre-briefing & post-briefing

· Always clarify EVA duration before it begins

· Devote more time to putting on radios & suits, check radios in advance,

· Respect communication protocols (not talk at the same time, only 1 person communicates with the Hab (this person must not be the driver)

· Hab must pay attention to coms with crew, speak slowly

· Keep close distance to each other & buddy system (always in pairs)

· Drone EVAs independent from other EVAs

· Backwards manoeuvres in rovers should be cleared

Destination: South of Poohs Corner

Coordinates (UTM NAD27 CONUS): Sector12S Northern 518781, Eastern 4250775

EVA Participants: Hector, Vittorio (Medic), Paolo and Zoe (Commander)

Road(s) and routes per MDRS Map: 1103 and Poohs Corner

Mode of travel: (walking or driving)? Driving [required for transportation of experiments]

Science Report – May 8th

Crew 212-LATAM-III

Submitted by Spacesuit Engineer: Camilo Andrés Zorro Mendoza

Spacesuit report

Sol: 3

Sim suit used: 8, 3, 9 and 5

Sim suit: 3
Voltage level before EVA: 12.9 V
Voltage level after EVA:
12.6 V

Sim suit: 5
Voltage level before EVA: 13.7 V
Voltage level after EVA:
12.9 V

Sim suit: 8
Voltage level before EVA: 12.8 V
Voltage level after EVA:
10.5 V

Sim suit: 9
Voltage level before EVA: 13.1 V
Voltage level after EVA:
12.7 V

General comments:

It was founded that four batteries of radios does not get charge. Also one sim suit recharger does not work. The radios was numbered.

Journalist Report – May 8th

[title Journalist Report – May 8th]

[category journalist-report]

Today we undertook an EVA despite the weather, however we had to stop after 1 hour. The EVA was to run three experiments simultaneously, soil collection for bacterial experiments in the science dome, first assembly of the drones and thermal experiments of the rover shielding. However, due to weather uncertainty and overconfidence with the suits none of the experiments could be completed on times.

Upon returning to the Hab, the problem solving experiment was conducted by all members of the crew. We then ate a broccoli and cheese soup with scones.

After lunch we worked on our individual experiments. The first set up of the flight profile for a microdrone using the radio was created for the drone projects and astronomy observations from the evening of the 6th were processed for galaxy NGC4236.

In addition interviews started for the blogs for the Engineer online blogs. These are part of a series of outreach tasks being undertaken.

Finally, success in the GreenHab resulted in a yield of carrots that have been harvested and will be used later this week for a meal.

Tonight is out cultural evening for the European members of our crew. They will be preparing food, music and games from their corresponding countries.

Copyright © The Mars Society. All rights reserved. | Main Site